OpenStudy (anonymous):

XYZ is a triangle with side lengths in the ratio 2 : 3 : 4. The perimeter of XYZ is 27 yards. Find the length of the shortest side.

OpenStudy (anonymous):



OpenStudy (skullpatrol):

The perimeter of a plane figure is the distance around it.

OpenStudy (skullpatrol):

@Megan_xoxo Does that hint help?

OpenStudy (anonymous):

no, im still confused.

OpenStudy (anonymous):

k (2 + 3 + 4) = 27 where k is a scaling factor. k=3 3*(2,3,4) = (6,9,12) or 6:9:12

OpenStudy (anonymous):

6 is the shortest side length.

OpenStudy (anonymous):

OH! okay, thank you!

OpenStudy (anonymous):

By the way, k = 3 is derived by solving the following for k: k (2 + 3 + 4) = 27 9k = 27 k = 3